OpenStudy (anonymous):

Can someone check my answers,, will medal ? What is the volume of a sphere that has a 14 m diameter? 205 m3 808 m3 1077 m3 1436 m3< I got this one .

OpenStudy (campbell_st):

well the radius is 7 m so the volume is \[V= \frac{4}{3} \times \pi \times 7^3\]

OpenStudy (campbell_st):

your answer is good. well done

OpenStudy (anonymous):

thank you okay ,, i got another one .

OpenStudy (anonymous):

If MS = 25 m, MT = 15 m, and TR = 5 m, what is the value of IS? 8.33 m 6.25 m< I got this one. 4 m 3 m

OpenStudy (campbell_st):

well that looks good as well

OpenStudy (anonymous):

. What is the measure of angle S? 12° 22° <I got this one 42° 44°

OpenStudy (campbell_st):

I would have said 42

OpenStudy (anonymous):

okay ,, could you explain the steps and formula to me ?

OpenStudy (campbell_st):

oops you're right... 1/2 the difference in the arc measures...

OpenStudy (anonymous):

Oh okay What is the approximate measure of angle R? 17.75° 29°< I got this one 43.6° 46.4°

OpenStudy (campbell_st):

well its a bit more than 45... since the triangle is almost isosceles

OpenStudy (anonymous):

I used the formula c^2 = a^2 + b^2 to get it . c^2 = 20^2 + 21^2 c^2 = 400 + 441 C^2 = 841 c^2 = 841 sq. c = 29

OpenStudy (campbell_st):

you can't use pythagoras' theorem to find an angle... you can use the tan ratio